the gcf of two numbers is 850 neither number is divisible by the other.what is the smallest that theese numbers could be
Answer by khwang(438)

If d is the gdf of A & B , then
A = d r and B = d s for some integers & s.
Also, if r & s (>1) are relatively prime then neither A nor B
is divisible by another.
Here,the gcf d= 850, choose the smallest relative prime pairs( >1) to
be 2 & 3.
That is,set A = 2*850 = 1700 and B = 3*850 = 2550
as the answer.
This question is not too hard, try to think about it.
Kenny